# Revealing Floating-Point Accumulation Orders in Software/Hardware Implementations

### v2

@@ Abstract
FPRev is a testing-based tool that non-intrusively reveals the accumulation order by analyzing the outputs of the tested implementation for distinct specially designed inputs. Employing FPRev, we showcase the accumulation orders of popular libraries (such as NumPy and PyTorch) on CPUs and GPUs (including GPUs with specialized matrix accelerators such as Tensor Cores). We also validate the efficiency of FPRev through extensive experiments. FPRev exhibits a lower time complexity compared to the basic solution. FPRev will be open-sourced on GitHub.
@@ 1 Introduction
Numerical reproducibility is critical in rigorous scenarios like finance [24, 23], where even minor inconsistencies in financial data are unacceptable. Software without verified numerical reproducibility is deemed risky or disqualified when applied to the rigorous scenarios. Unfortunately, with the rapid evolution of heterogeneous hardware and the fast iteration of diverse software stacks, reproducibility of AccumOps has become increasingly challenging. Existing implementations rarely disclose their accumulation orders, hindering reproducible AccumOp development.
@@ 2.1.1 On canonical CPUs and GPUs
Second, the same software may behave differently across different hardware. Different CPUs and GPUs vary in architecture, number of cores, SIMD width, cache size, etc. Consequently, for performance optimization, software may adjust the accumulation order based on the specific hardware characteristic. Although order-independent algorithms have been proposed [5, 6, 4], which ensure consistent results regardless of the accumulation order, they are highly inefficient and thus rarely used in industry.
@@ 3.1 Motivation
An example application of the tool is in developing reproducible AccumOps, which are key to software reproducibility and stability [10, 27]. Developers must ensure that the accumulation order remains consistent across systems to maintain the reproducibility of AccumOps. If developers can determine the accumulation order, they can use it as a specification to guide their development process. In addition, when porting software to a new system, developers need a rigorous way to verify the equivalence of AccumOps between two systems. This can be achieved by comparing the accumulation orders of the AccumOps implemented on two systems.
@@ 3.2 Problem definition
We formulate the summation operation as follows. The floating-point addition is performed $n-1$ times in a predetermined order to calculate the sum of $n$ floating-point numbers. We assume that the accumulation order is unknown but is uniquely determined by the given implementation on specific hardware. Therefore, randomized implementations and those where the order depends on the values of the summands are out of scope.
@@ 4.2 Step 2: analyzing the accumulation order from the outputs
Take Algorithm 1 as an example SumImpl, whose accumulation order is depicted in Figure 2. If $i=2$ and $j=4$, then the array $A^{2,4}$ is $(1,1,M,1,-M,1,1,1)$. Computing the sum of $A^{2,4}$ with the example SumImpl, the 3rd summand and the intermediate sum of the 0th and 1st summands are masked by $M$ (the 2nd summand); the 5th summand is masked by $-M$ (the 4th summand). Therefore, in total, $n_{\mathrm{masked}}^{2,4}=4$. In contrast, the 6th and 7th summands and their intermediate sum are not added to $M$ or $-M$, so $n_{\mathrm{not\ masked}}^{2,4}=\textsc{SumImpl}(A^{2,4})=2$.
@@ 5.3 Time complexity and correctness
The time complexity of FPRev is $O(n^{2}t(n))$ and $\Omega(nt(n))$, following the same analysis in Section 5.1.3. The correctness of it is also guaranteed by design and can be proven following the same process in Section 4.4.
@@ 6.2 PyTorch’s implementation on GPUs
Figure 4: The accumulation orders of PyTorch’s $32\times 32\times 32$ matrix multiplication on Tensor Cores.
@@ 7.1 Experiment design
Figure 5: Run time of applying NaiveSol, BasicFPRev, and FPRev to the summation functions in NumPy, PyTorch, and JAX. The vertical axis represents run time in seconds. The horizontal axis represents the number of summands $n$.
Figure 6: Run time of applying BasicFPRev and FPRev to the dor product, matrix-vector multiplication, and matrix multiplication functions in NumPy. The vertical axis represents run time in seconds. The horizontal axis represents the number of summands $n$.
@@
To answer the RQs, we measure the run time (wall-clock time) of applying our solutions to tested libraries and operations. We implement FPRev (Algorithm 4) in Python (version 3.11). For comparison, we also implement the basic solution (Algorithm 2), denoted by BasicFPRev.
@@ 7.2 RQ1: How efficient is FPRev when applied to different libraries?
Figure 7: Run time of applying BasicFPRev and FPRev to the matrix multiplication functions in PyTorch on different CPUs and GPUs. The vertical axis represents run time in seconds. The horizontal axis represents the number of summands $n$.
@@
We begin with the number of summands $n=4$, and increment $n$ until the run time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Figure 5. The red curves indicate that the run time of NaiveSol grows exponentially as $n$ grows. The results substantiate the $O(4^{n}/n^{3/2}\cdot t(n))$ time complexity of NaiveSol. The green and blue lines show that the run time of BasicFPRev and FPRev grows polynomially. The different slopes also demonstrate that the run time of BasicFPRev is longer than that of FPRev, and increases more rapidly as $n$ increases. This is because the time complexity of BasicFPRev is $\Theta(n^{2}t(n))$, while that of FPRev is $\Omega(nt(n))$ and $O(n^{2}t(n))$.
@@ 7.3 RQ2: How efficient is FPRev when applied to different operations?
Similarly, we begin with $n=4$, and increment $n$ until the run time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Table 6. The different slopes indicate that the time complexity of BasicFPRev is higher than that of FPRev. In addition, as the time complexity of the workload increases, the growth speed of the runtime with regard to $n$ accelerates. Therefore, the speedup of FPRev over BasicFPRev is more pronounced as the workload is more complex. For example, for $n=256$, FPRev is 13.0×13.0$\times$13.0 × as fast as BasicFPRev for dot product, 32.3×32.3$\times$32.3 × for matrix-vector multiplication, and 82.1×82.1$\times$82.1 × for matrix multiplication.
@@ 7.4 RQ3: How efficient is FPRev on different CPUs and GPUs?
For RQ3, we test the single-precision matrix multiplication in PyTorch on the CPUs and GPUs listed in Section 6. Similarly, we begin with $n=4$, and increment $n$ until the run time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Figure 7. The results demonstrate consistent improvements in the runtime of FPRev compared to BasicFPRev. Therefore, FPRev is consistently more efficient than BasicFPRev on different devices.
@@ 8 Discussion
8.1 Threats to validity
Internal. The precision of the floating-point data type can limit the validity of FPRev. For float32, which has a precision of 24 bits, the maximum number of summands ($n$) that FPRev supports is $167772172^{24}+1=$ 16777217. For larger numbers, the sum of $n-2$ ones cannot be represented precisely by float32, so the sum of masked all-one arrays may be incorrect. This problem could be solved by dynamically replacing the multiple ones corresponding to a constructed subtree with their sum and multiple zeros. This solution does not affect efficiency.
@@ 8.1 Threats to validity
For low-precision data types, there is an additional issue of the floating-point swamping because $M\ll n$ may not hold. For example, the maximum exponent of float16 is 15. If $M=2^{15}$ and $\mu=256$, ±$\pm M+\mu\neq\pm M$. To solve this problem, we can replace the ones in the masked all-one arrays with smaller numbers (e.g., $2^{-24}$), and scale the sum back to an integer between 0 and $n-2$ when calculating $l^{i,j}$. This solution does not affect efficiency, either.
External. If the tested implementation has functional defects, FPRev may not yield the correct result. Nonetheless, we can still extract some diagnostic information from the abnormal output of FPRev. For example, if a conversion from float64 to float32 occurs during summation, FPRev will detect overflows, indicating the presence of the conversion.
8.2 Extensibility
@@
Section 6 has demonstrated that FPRev can be applied to most popular numerical libraries. FPRev can be applied to other implementations as long as they fall within the scope detailed in Section 3. FPRev will also work for collective communication primitives if their accumulation order is predetermined. To further extend our tool to other functions based on special summation algorithms, the algorithms must satisfy the property $l^{i,j}=n-\textsc{SumImpl}(A^{i,j})$. Thus, our methods can be abstracted into a template to reveal the accumulation order.
@@ 9 Conclusion
In this paper, we introduce FPRev, a diagnostic tool for revealing the accumulation order in software and hardware implementations through numerical testing. It can help verify and facilitate the development of reproducible software. As a case study, FPRev reveal the undisclosed accumulation orders of prevalent numerical libraries such as NumPy and PyTorch on different CPUs and GPUs. We also demonstrated the efficiency FPRev through experiments covering various implementations and devices. Our source code will be made available after the double-blind review process, encouraging further investigation and improvement by the research community.

### v3

@@ Abstract
FPRev is a testing-based tool that non-intrusively reveals the accumulation order by analyzing the outputs of the tested implementation for distinct specially designed inputs. Employing FPRev, we showcase the accumulation orders of popular libraries (such as NumPy and PyTorch) on CPUs and GPUs (including GPUs with specialized matrix accelerators such as Tensor Cores). We also validate the efficiency of FPRev through extensive experiments. FPRev exhibits a lower time complexity compared to the basic solution. FPRev is open-sourced at https://github.com/peichenxie/FPRev.
@@ 1 Introduction
Numerical reproducibility is critical in scientific computing [12, 33, 28], high-performance computing [35], database management system [22], deep learning [31, 4], etc. Particularly, software without verified numerical reproducibility is deemed risky or disqualified when applied to safety-critical or rigorous scenarios like aerospace or banking, where even minor inconsistencies in data are unacceptable. Unfortunately, with the rapid evolution of heterogeneous hardware and the fast iteration of diverse software stacks, reproducibility of AccumOps has become increasingly challenging. Existing implementations rarely disclose their accumulation orders, hindering reproducible AccumOp development.
@@ 2.1.1 On canonical CPUs and GPUs
Second, the same software may behave differently across different hardware. Different CPUs and GPUs vary in architecture, number of cores, SIMD width, cache size, etc. Consequently, for performance optimization, software may adjust the accumulation order based on the specific hardware characteristic. Specifically, library developers implement various techniques (e.g., different configurations of loop unrolling, block partitioning, cache optimization, and vectorization) for performance tuning, resulting in different accumulation orders. Additionally, auto-tuners (e.g., Triton [34] and TVM [5]) are often used to search for optimal configurations, given the complexity of performance factors such as instruction pipelining and dynamic frequency scaling.
Although order-independent algorithms have been proposed [7, 8, 6], which ensure consistent results regardless of the accumulation order, they are highly inefficient and thus rarely used in industry.
@@ 3.1 Motivation
An example application of the tool is in developing reproducible AccumOps, which are key to software reproducibility and service consistency [12, 33, 25]. Developers must ensure that the accumulation order remains consistent across systems to maintain the reproducibility of AccumOps. If developers can determine the accumulation order, they can use it as a specification to guide their development process. In addition, when porting software to a new system, developers need a rigorous way to verify the equivalence of AccumOps between two systems. This can be achieved by comparing the accumulation orders of the AccumOps implemented on two systems.
@@ 3.2 Problem definition
We formulate the summation operation as follows. The floating-point addition is performed $n-1$ times in a predetermined order to calculate the sum of $n$ floating-point numbers. We assume that the accumulation order is unknown but is uniquely determined by the given implementation on specific hardware. Therefore, randomized implementations and those where the order depends on the values of the summands are out of scope [^1] While randomized or input-dependent summation algorithms are possible, they tend to be inefficient. Similarly, algorithms using AtomicAdd are affected by thread scheduling, but we did not find popular libraries implementing such algorithms..
@@ 4.2 Step 2: analyzing the accumulation order from the outputs
Take Algorithm 1 as an example SumImpl, whose accumulation order is depicted in Figure 2.222We compile it with different compiler versions and optimization flags, and note that they do not change the accumulation order, even if they might in theory. We suspect such changes can be found on more complex math expressions instead of summations. If $i=2$ and $j=4$, then the array $A^{2,4}$ is $(1,1,M,1,-M,1,1,1)$. Computing the sum of $A^{2,4}$ with the example SumImpl, the 3rd summand and the intermediate sum of the 0th and 1st summands are masked by $M$ (the 2nd summand); the 5th summand is masked by $-M$ (the 4th summand). Therefore, in total, $n_{\mathrm{masked}}^{2,4}=4$. In contrast, the 6th and 7th summands and their intermediate sum are not added to $M$ or $-M$, so $n_{\mathrm{not\ masked}}^{2,4}=\textsc{SumImpl}(A^{2,4})=2$.
@@ 5.3 Time complexity and correctness
Following the same analysis in Section 5.1.3, the time complexity of FPRev is $O(n^{2}t(n))$ and $\Omega(nt(n))$, and the probability of the worst-case time complexity $O(n^{2}t(n))$ is low. The correctness of it is also guaranteed by design and can be proven following the same process in Section 4.4.
@@ 6.2 PyTorch’s implementation on GPUs
Figure 4: The accumulation orders of PyTorch’s half-precision $32\times 32\times 32$ matrix multiplication on Tensor Cores.
@@ 7.1 Experiment design
Figure 5: Execution time of applying NaiveSol, BasicFPRev, and FPRev to the summation functions in NumPy, PyTorch, and JAX. The vertical axis represents execution time in seconds. The horizontal axis represents the number of summands $n$.
Figure 6: Execution time of applying BasicFPRev and FPRev to the dor product, matrix-vector multiplication, and matrix multiplication functions in NumPy. The vertical axis represents execution time in seconds. The horizontal axis represents the number of summands $n$.
@@
To answer the RQs, we measure the execution time (wall-clock time) of applying our solutions to tested libraries and operations. We implement FPRev (Algorithm 4) in Python (version 3.11). For comparison, we also implement the basic solution (Algorithm 2), denoted by BasicFPRev.
@@ 7.2 RQ1: How efficient is FPRev when applied to different libraries?
Figure 7: Execution time of applying BasicFPRev and FPRev to the matrix multiplication functions in PyTorch on different CPUs and GPUs. The vertical axis represents execution time in seconds. The horizontal axis represents the number of summands $n$.
@@
We begin with the number of summands $n=4$, and increment $n$ until the execution time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Figure 5. The red curves indicate that the execution time of NaiveSol grows exponentially as $n$ grows. The results substantiate the $O(4^{n}/n^{3/2}\cdot t(n))$ time complexity of NaiveSol. The green and blue lines show that the execution time of BasicFPRev and FPRev grows polynomially. The different slopes also demonstrate that the execution time of BasicFPRev is longer than that of FPRev, and increases more rapidly as $n$ increases. This is because the time complexity of BasicFPRev is $\Theta(n^{2}t(n))$, while that of FPRev is $\Omega(nt(n))$ and $O(n^{2}t(n))$.
@@ 7.3 RQ2: How efficient is FPRev when applied to different operations?
Similarly, we begin with $n=4$, and increment $n$ until the execution time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Table 6. The different slopes indicate that the time complexity of BasicFPRev is higher than that of FPRev. In addition, as the time complexity of the workload increases, the growth speed of the runtime with regard to $n$ accelerates. Therefore, the speedup of FPRev over BasicFPRev is more pronounced as the workload is more complex. For example, for $n=256$, FPRev is 13.0×13.0$\times$13.0 × as fast as BasicFPRev for dot product, 32.3×32.3$\times$32.3 × for matrix-vector multiplication, and 82.1×82.1$\times$82.1 × for matrix multiplication.
@@ 7.4 RQ3: How efficient is FPRev on different CPUs and GPUs?
For RQ3, we test the single-precision matrix multiplication in PyTorch on the CPUs and GPUs listed in Section 6. Similarly, we begin with $n=4$, and increment $n$ until the execution time exceeds one second. Each experiment is carried out 10 times, and the arithmetic mean of the 10 results is reported in Figure 7. The results demonstrate consistent improvements in the runtime of FPRev compared to BasicFPRev. Therefore, FPRev is consistently more efficient than BasicFPRev on different devices.
@@ 8 Discussion
8.1 Limitation and mitigation
8.1.1 Dynamic range of the input data type
@@
When applying FPRev to data types with low dynamic range, the mask $M$ may be too small to effectively mask the sum of ones. For example, the maximum value of the 8-bit floating-point number with 4 exponent bits (FP8-e4m3) defined in [21] is $1.75\times 2^{8}$, so the condition $M\gg n$ may not hold, and ±$\pm M+\sigma\neq\pm M$ for $20\leq\sigma\leq n-$ 2. To mitigate the issue, we can replace the ones in the masked all-one arrays with smaller numbers (e.g., $2^{-9}\times 2^{-9}$ for FP8-e4m3 matrix multiplication), and scale the sum back to an integer between 0 and $n-2$ when calculating $l^{i,j}$. This solution does not affect efficiency.
8.1.2 Precision of the accumulator
The precision of the floating-point accumulator can limit the input size that FPRev supports. For example, float32 has a precision of 24 bits, so the maximum number of summands ($n$) that FPRev supports is $167772172^{24}+1=$ 16777217 for float32 accumulation operations. For larger numbers, the sum of $n-2$ ones cannot be represented precisely in float32, so the sum of masked all-one arrays may be incorrect. This issue can be mitigated by dynamically replacing the multiple ones corresponding to a constructed subtree with one and multiple zeros, as if compressing the constructed subtree into one node.
Specifically in the BuildSubtree($I$) function of FPRev (Algorithm 4), the computation of $\textsc{sum}(A^{i,j})=0$ is accurate for $j$s such that $l^{i,j}=n$, so we can build the subtree for them in the end. We extract those $j$s to $J=\{j:l^{i,j}=n\}$ after computing $L_{i}=\{l^{i,j}:j\in I-\{i\}\}$. Then, we set the values at $J$ to 0, and build the subtree for $I-J$ recursively (which results in a smaller subproblem). After the subtree for $I-J$ is constructed, we set the values at $J$ back to 1.0, and set the values at $I-J-\{i\}$ to 0. Now, the constructed subtree (containing $I-J$) is treated as a node, represented by node #$\#i$. Next, we run the original tree construction algorithm (the last iteration of the for-loop, i.e., $l=|All|$) for $J$, and then the whole tree is constructed.
Combining the two mitigation techniques, the modified version of FPRev is shown in Algorithm 5. This version is applicable to data types with low dynamic range and low accumulation precision, such as 16-bit and 8-bit floating-point formats (including BF16, FP16, FP8-e5m2, and FP8-e4m3 on recent Tensor Cores).
Algorithm 5 The modified version of FPRev.
Implementation SumImpl, the number of summands $n$, large value $M$, and tiny value $e$
Summation tree of SumImpl
function ModifiedFPRev($\textsc{SumImpl},n,M,e$)
     function BuildSubtree($I,All$)
         $T\leftarrow\varnothing$
         if $|I|=1$ then $\triangleright$ stop condition
              return $T$
         $i\leftarrow\min(I)$; $L_{i}\leftarrow\varnothing$
         for $j\in I-\{i\}$ do $\triangleright$ calculate $l^{i,j}$ on demand
              $A^{i,j}_{k}\leftarrow e$ for $k\in All$; $A^{i,j}_{i}\leftarrow M$; $A^{i,j}_{j}\leftarrow-M$
              $l^{i,j}\leftarrow|All|-\textsc{SumImpl}(A^{i,j})/e$
              $L_{i}\leftarrow L_{i}\cup\{l^{i,j}\}$
         $J\leftarrow\{j:l^{i,j}=\max(L_{i})\}$
         $A^{i,j}_{k}\leftarrow 0$ for $k\in J$triangleright Ignoring $J$
         $(T,n_{\mathrm{leaves}}^{T_{c}})\leftarrow\textsc{BuildSubtree}(I-J,All-J)$
         $r\leftarrow\mathrm{GetRoot}(T)$
         $A^{i,j}_{k}\leftarrow e$ for $k\in J$
         $K\leftarrow I-J-\{i\}$
         $A^{i,j}_{k}\leftarrow 0$ for $k\in K$triangleright Treating $I-J$ as $\{i\}$
         $(T^{\prime},n_{\mathrm{leaves}}^{T_{c}})\leftarrow\textsc{BuildSubtree}(J,All-K)$
         $T\leftarrow T\cup T^{\prime}$
         if $|J|=n_{\mathrm{leaves}}^{T_{c}}$ then $\triangleright$ $T^{\prime}=T_{c}$
              $T\leftarrow T\cup\{(r,r+n),(\mathrm{GetRoot}(T^{\prime}),r+n)\}$
         $\triangleright$ $T^{\prime}\subset T_{c}$
              $T\leftarrow T\cup\{(r,\mathrm{GetRoot}(T^{\prime}))\}$
         return $(T,\max(L_{i}))$
     $All\leftarrow\{0,1,...,n-1\}$
     $(T,n_{\mathrm{leaves}}^{T_{c}})\leftarrow\textsc{BuildSubtree}(All,All)$
     return $T$
8.2 Extensibility and future work
@@
Section 6 has demonstrated that FPRev can be applied to popular numerical libraries like NumPy and PyTorch. FPRev can be applied to other accumulation implementations as long as they fall within the scope detailed in Section 3. In practice, we find most popular libraries have deterministic reduction orders and fall into the scope. FPRev also works for accumulation operations in collective communication primitives, such as the AllReduce operation, if their accumulation order is predetermined.
To further extend our tool to other functions based on special summation algorithms, the algorithms must satisfy the property $l^{i,j}=n-\textsc{SumImpl}(A^{i,j})$ or its variant in Algorithm 5. For example, the next generation of Tensor Core will support the microscaling data format [30], including the 4-bit and 6-bit formats MXFP4 and MXFP6. If their dynamic range and accumulator precision permit and the property holds, our methods can reveal the accumulation order within a block of microscaling numbers. Then, we can treat a block as one summand, and use FPRev to construct the summation tree for the summation of the blocks, and then expand each block to a subtree.
In addition to revealing accumulation orders, we plan to extend our methods to detect more floating-point behaviors in matrix accelerators. For example, we can determine the rounding mode and the precision of the accumulator of Tensor Cores by enumerating $n=1,2,...$ and checking the result of $2^{n}+1.75-2^{n}$. We are designing more numerical experiments to identify how the block fused multiply-add is conducted. Then, with the information detected, we can model the exact behavior of the hardware matrix accelerators.
Another direction is further optimizing the efficiency of FPRev. For example, we can randomize the selection of $i\in I$ in the FPRev algorithm, as if selecting the random pivot in quick sort. This might reduce the expected time complexity of FPRev.
@@ 9 Conclusion
In this paper, we introduce FPRev, a diagnostic tool for revealing the accumulation order in software and hardware implementations through numerical testing. It can help verify and facilitate the development of reproducible software. As a case study, FPRev reveal the undisclosed accumulation orders of prevalent numerical libraries such as NumPy and PyTorch on different CPUs and GPUs. We also demonstrated the efficiency FPRev through experiments covering various implementations and devices. Our source code is available at https://github.com/peichenxie/FPRev, encouraging further investigation and improvement by the research community.
Appendix A Artifact Appendix
Abstract
The repository includes the source code of FPRev and the source code for reproducing the experiments of the paper. The following content includes main claims that can be verified via experiments, and instructions to reproduce the experiments.
Scope
The main claims include:
  1. 1.
    FPRev is functional to reveal the floating-point accumulation orders in common implementations. This claim is detailed by Section 6 “Case study” in the paper. To verify this claim, run python experiments/casestudy.py and check the output files in the outputs directory.
    1. (a)
      outputs/Numpy*.pdf represents the revealed accumulation orders for NumPy, as discussed in Section 6.1 “NumPy’s implementation on CPUs”. Among them, outputs/NumpyGEMV8.pdf corresponds to Figure 3 of the paper, if the CPU models are consistent to those in the paper.
    2. (b)
      outputs/Torch*.pdf represents the revealed accumulation orders for PyTorch, as discussed in Section 6.2 “PyTorch’s implementation on GPUs”. Among them, outputs/TorchF16GEMM32.pdf corresponds to Figure 4 of the paper, if the GPU models are consistent to those in the paper.
  2. 2.
    FPRev is efficient. This claim is detailed by Section 7 “Performance evaluation” in the paper. To verify this claim, run python experiments/rq1.py, python experiments/rq2.py, and python experiments/rq3.py, and check the output files in the outputs directory.
    1. (a)
      outputs/rq1.csv provides the results of Section 7.2 “RQ1: How efficient is FPRev when applied to different libraries?”. It corresponds to Figure 5 if the CPU model is consistenst to that in the paper.
    2. (b)
      outputs/rq2.csv provides the results of Section 7.3 “RQ2: How efficient is FPRev when applied to different operations?”. It corresponds to Figure 6 if the CPU model is consistenst to that in the paper.
    3. (c)
      outputs/rq3.csv provides the results of Section 7.4 “RQ3: How efficient is FPRev on different CPUs and GPUs?”. It corresponds to Figure 7 if the CPU and GPU models are consistenst to those in the paper.
Contents
Installation
sudo apt install graphviz
git clone https://github.com/peichenxie/FPRev.git
cd FPRev
pip install .
pip install -r experiments/requirements.txt
Running experiments
  • To reproduce the results in Section 6 (Case study), run python experiments/casestudy.py on different hardware models.
  • To reproduce the results in Section 7.2 (RQ1: How efficient is FPRev when applied to different libraries?), run python experiments/rq1.py.
  • To reproduce the results in Section 7.3 (RQ2: How efficient is FPRev when applied to different operations?), run python experiments/rq2.py.
  • To reproduce the results in Section 7.4 (RQ3: How efficient is FPRev on different CPUs and GPUs?), run python experiments/rq3.py on different hardware models.
Then, check the output files in the outputs directory. See outputs/README.md for more information.
Hosting
Requirements
The artifact requires the following platform:
  • GPU: NVIDIA V100 or newer
  • OS: Ubuntu 22.04
  • Software: Python (version 3.11)
If you wish to reproduce the results in the paper, please use the identical CPU and GPU models:
  1. 1.
    CPU: Intel Xeon E5-2690 v4 (24 v-cores), GPU: NVIDIA V100 (5120 CUDA cores)
  2. 2.
    CPU: AMD EPYC 7V13 (24 v-cores), GPU: NVIDIA A100 (6912 CUDA cores)
  3. 3.
    CPU: Intel Xeon Silver 4210 (40 v-cores), GPU: NVIDIA H100 (16896 CUDA cores)